OpenStudy (anonymous):

Please Helppp !!!! Solve 5x2 = -30x - 65. x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i

OpenStudy (anonymous):

Do you know the Quadratic formulae?

OpenStudy (anonymous):

yes, i do know the quadratic formula

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac}}{ 2a}\]

OpenStudy (anonymous):

Rearrage your eqn in quadratic form \[ax^2 + bx + c = 0\]subst a, b, c and solve for x

OpenStudy (anonymous):

hold on , let me try

OpenStudy (anonymous):

i'm kind of confused

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

Putting all together : \[\text{Putting all together: } x = \frac{ -30\pm \sqrt{30^2 - 4*5*65}}{ 2*5}\]

OpenStudy (anonymous):

\[ax ^{2}+bx+c+0 \] When i substitute them i get \[5x ^{2}+30x-+65=0\] ?

OpenStudy (anonymous):

When you move the expression on the right to the left, the -ve signs turn +ve. You remember?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

\[\text{So you get : } 5x ^{2}+30x +65=0\]

OpenStudy (anonymous):

i also remember that when solving the quadratic formula there was something that i have to do first, and i can't remember if it's the bottom half of the equation that i start with, or the part that says 30^2 - 4*5*65

OpenStudy (anonymous):

\[\text{Simplify this further: } x = \frac{ -30\pm \sqrt{30^2 - 4*5*65}}{ 2*5} \rightarrow x = \frac{ -30\pm \sqrt{900 - 1300}}{10}\]

OpenStudy (anonymous):

Does that make sense now?

OpenStudy (anonymous):

\[x = \frac{ -30\pm \sqrt{900 - 1300}}{10} = \frac{ -30 }{ 10 } \pm \frac{ \sqrt{-400} }{ 10 }\]

OpenStudy (anonymous):

and next you do 900 - 1300 and get \[x= -30\pm \sqrt{-400} / 10\]

OpenStudy (anonymous):

or do i make the 400 positive ?

OpenStudy (anonymous):

Nope, that's where the imaginary number 'i' comes in, since we do not the negative square root of a number...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have it rather like this\[x = -3 \pm \frac{ \sqrt{400i} }{ 10 } \rightarrow x = -3 \pm \frac{ 20i }{ 10 } \]

OpenStudy (anonymous):

\[x = -3 \pm 2i\]

OpenStudy (anonymous):

wow that is confusing, haha

OpenStudy (anonymous):

It's very important you understand every bit of the process that got us here. So, what confusion can we clear up now?

OpenStudy (anonymous):

the imaginary number is whats confusing me

OpenStudy (anonymous):

Yeah, it's just a rule. You know that now, it won't confuse you any more :)

OpenStudy (anonymous):

ohh i see, the 0's cancel out

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

And you notice that we took the sqrt of 400 too

OpenStudy (anonymous):

wait, what happened with the sort of 400 ? it canceled out too ?

OpenStudy (anonymous):

You just keep the i out of it and take the square root of its coefficient.

OpenStudy (anonymous):